OpenStudy (anonymous):

HELP PLEASE Find f(x) and g(x), so that the function can be described as y=f(g(x)). y=|8x+9| How would you do this?????!

OpenStudy (anonymous):

Do you understand what y=f(g(x)) is asking you to do?

OpenStudy (anonymous):

It shows a composition of two functions. The function g(x) is being taken as the input to the function f(x)

OpenStudy (anonymous):

f(x) = x + 9, g(x) = 4/x^2

OpenStudy (anonymous):

f(x) = |x|, g(x) = 8x+9

OpenStudy (anonymous):

so y = |8x+9| = f(g(x))

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you both so much!